OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer? (1 point)

OpenStudy (texaschic101):

x = 1st integer x + 1 = 2nd integer x + 2 = 3rd integer (x) + (x + 1) + (x + 2) = 267 -- combine like terms 3x + 3 = 267 -- subtract 3 from both sides 3x = 267 - 3 3x = 264 -- divide by 3 x = 264/3 = x = 88 ==>1st integer x + 1 = 88 + 1 = 89 ==>2nd integer x + 2 = 88 + 2 = 90 ==> 3rd integer check.. 88 + 89 + 90 = 267 267 = 267 (correct) Largest number is 90.